OpenStudy (lgbasallote):

OpenStudy user A joined OpenStudy 6 months before OpenStudy user B. After how many months will OpenStudy user B's time spent in OpenStudy be half of OpenStudy user A's time spent in OpenStudy?

Parth (parthkohli):

Take user A as \(x + 6\), and user B as \(x\).

OpenStudy (lgbasallote):

so then it will be x + m = (x + 6 + m)/2?

Parth (parthkohli):

Though I did this through trial and error and got 6 months as the answer =))

OpenStudy (lgbasallote):

where m is months later of course

OpenStudy (lgbasallote):

lol what's the algebraic solution?

Parth (parthkohli):

Thinking :p

OpenStudy (lgbasallote):

i think my equation makes sense :/

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2x = (x + 6) + 6 }\) Aha

Parth (parthkohli):

Solve for \(x\).

Parth (parthkohli):

\(2x = x + 12\) \( \color{Black}{\Rightarrow x = 12 }\) wait :/

OpenStudy (lgbasallote):

why +6?

Parth (parthkohli):

Well, if we say that the answer is 'after x months', then the time spent for the newbie would be x months and the other one x + 6 months

OpenStudy (lgbasallote):

2(x+m) = x + 6 + m x + m = 6 therefore 6 months?

Parth (parthkohli):

Oh yeah we could make a system :P

Parth (parthkohli):

Correct

OpenStudy (lgbasallote):

thanks

Parth (parthkohli):

Lol you did the problem on your own =))

OpenStudy (lgbasallote):

a follow up question was asking what if A came 11 months before B?

Parth (parthkohli):

Where are these questions, seriously :P

OpenStudy (lgbasallote):

the answer's not in my book =_=

OpenStudy (lgbasallote):

my book?

Parth (parthkohli):

Your book knows OS lol

OpenStudy (lgbasallote):

yup :p i wrote it ^_^

Parth (parthkohli):

lol

OpenStudy (lgbasallote):

so what if 11 months?

Parth (parthkohli):

\( \color{Black}{\Rightarrow x + m = 11 }\) \( \color{Black}{\Rightarrow 2(x + m) = x +11 + m }\)

OpenStudy (lgbasallote):

so x + m = 11...

OpenStudy (lgbasallote):

that's weird..

Parth (parthkohli):

\( \color{Black}{\Rightarrow 22 = 11 + 11 }\) \( \color{Black}{\Rightarrow 22 = 22 }\) :/

OpenStudy (lgbasallote):

the time difference between them is the amount of time it takeS B to be half of A?

Parth (parthkohli):

lol I'm noticing something here

Parth (parthkohli):

If x + m = 6, then it takes 6 months If x + m = 11, then it takes 11 months

Parth (parthkohli):

I am a magician =.=

OpenStudy (lgbasallote):

something is definitely weird >.<

OpenStudy (lgbasallote):

let's say i joined July and you joined april...that means 9 months after april i should be double of you.. oh yeah makes sense lol

Parth (parthkohli):

lol that's easy xP

Parth (parthkohli):

See what we proved with trial and error mwahahaha

OpenStudy (lgbasallote):

lol

Parth (parthkohli):

TRIAL AND ERROR FTW